OpenStudy (anonymous):

Which measurement is a possible side length for a triangle with two sides measuring 9cm and 4cm? A) 5cm B) 6cm C) 3cm D)4cm Can someone explain how exactly to do this?

OpenStudy (anonymous):

Use the Pythagorean equation.

OpenStudy (anonymous):

A^2 + B^2 = C^2

OpenStudy (anonymous):



OpenStudy (anonymous):

there is a rule for triangles...the sum of two sides must be greater than the third side A) 5...suppose i take 5 and 4 as two sides...this must be greater than 9..but it's not..therefore it's not 5 let's look at D) 4 4+4 is not greater than 9 either so that's not it continue with this logic...do you know the answer?

OpenStudy (anonymous):

So it would be 6 ... Because 6+4=10 and 10 is greater than 9

OpenStudy (anonymous):

you got it!

OpenStudy (anonymous):

6 + 9 is also greater than 4 which further supports your answer

OpenStudy (anonymous):

Awesome thanks :)